# Measuring and Controlling Solution Degeneracy across Task-Trained Recurrent Neural Networks

### v2

@@ Abstract
Task-trained recurrent neural networks (RNNs) are widely used in neuroscience and machine learning to model dynamical computations. To gain mechanistic insight into how neural systems solve tasks, prior work often reverse-engineers individual trained networks. However, different RNNs trained on the same task and achieving similar performance can exhibit strikingly different internal solutions—a phenomenon known as solution degeneracy. Here, we develop a unified framework to systematically quantify and control solution degeneracy across three levels: behavior, neural dynamics, and weight space. We apply this framework to 3,400 RNNs trained on four neuroscience-relevant tasks—flip-flop memory, sine wave generation, delayed discrimination, and path integration—while systematically varying task complexity, learning regime, network size, and regularization. We find that higher task complexity and stronger feature learning reduce degeneracy in neural dynamics but increase it in weight space, with mixed effects on behavior. In contrast, larger networks and structural regularization reduce degeneracy at all three levels. These findings empirically validate the Contravariance Principle and provide practical guidance for researchers aiming to tailor RNN solutions—whether to uncover shared neural mechanisms or to model individual variability observed in biological systems. This work provides a principled framework for quantifying and controlling solution degeneracy in task-trained RNNs, offering new tools for building more interpretable and biologically grounded models of neural computation.
@@ 1 Introduction
Recurrent neural networks (RNNs) are widely used in machine learning and computational neuroscience to model dynamical processes. They are typically trained with standard nonconvex optimization methods and have proven useful as surrogate models for generating hypotheses about the neural mechanisms underlying task performance [76, 72, 5, 57, 79, 18]. Traditionally, the study of task-trained RNNs has focused on reverse-engineering a single trained model, implicitly assuming that networks trained on the same task would converge to similar solutions—even when initialized or trained differently. However, recent work has shown that this assumption does not hold universally, and the solution space of task-trained RNNs can be highly degenerate: networks may achieve the same level of training loss, yet differ in out-of-distribution (OOD) behavior, internal representations, neural dynamics, and connectivity [78, 26, 30, 27, 58]. For instance, [56] found that while trained RNNs may share certain topological features, their representational geometry can vary widely. Similarly, [78] showed that task-trained networks can develop qualitatively distinct neural dynamics and OOD generalization behaviors.
These findings raise fundamental questions about the solution space of task-trained RNNs: What factors govern the solution degeneracy across independently trained RNNs? When the solution space of task-trained RNNs is highly degenerate, to what extent can we trust conclusions drawn from a single model instance? While feedforward networks have been extensively studied in terms of how weight initialization and stochastic training (e.g., mini-batch gradients) lead to divergent solutions, RNNs still lack a systematic and unified understanding of the factors that govern solution degeneracy [16, 77, 15, 31, 23, 34, 49, 40, 12, 24, 45].
Cao and Yamins [10] proposed the Contravariance Principle, which posits that as the computational objective (i.e., the task) becomes more complex, the solution space should become less dispersed—since fewer models can simultaneously satisfy the stricter constraints imposed by harder tasks. While this principle is intuitive and compelling, it has thus far remained largely theoretical and has not been directly validated through empirical studies.
@@
In this paper, we introduce a unified framework for quantifying solution degeneracy at three levels: behavior, neural dynamics, and weight space. As illustrated in Figure 1 , we quantify degeneracy across behavior, dynamics, and weights, and examine how it is shaped by four key factors. Leveraging this framework, we isolate four key factors that control solution degeneracy—task complexity, learning regime, network width, and structural regularization. We apply this framework in a large-scale experiment, training 50 independently initialized RNNs on each of four neuroscience-relevant tasks. By systematically varying task complexity, learning regime, network width, and regularization, we map how each factor shapes degeneracy across behavior, dynamics, and weights. We find that as task complexity increases—whether via more input–output channels, higher memory demand, or auxiliary objectives—or as networks undergo stronger feature learning, their neural dynamics become more consistent, while their weight configurations grow more variable. In contrast, increasing network size or imposing structural regularization during training reduces variability at both the dynamics and weight levels. At the behavioral level, each of these factors reliably modulates behavioral degeneracy; however, the relationship between behavioral and dynamical degeneracy is not always consistent.
Table 1 summarizes how task complexity, learning regime, network size, and regularization affect degeneracy across levels. In both machine learning and neuroscience, the optimal level of degeneracy may vary depending on the specific research questions being investigated. This framework offers practical guidance for tailoring training to a given goal—whether encouraging consistency across models [42], or promoting diversity across learned solutions [52, 21, 69].
@@
  • A systematic sweep of four factors—task complexity, feature learning, network size, and regularization—and a summary of their effects across levels (Table 1), with practical guidance for tuning consistency vs. diversity [42, 52, 21, 69].
@@
The remainder of the paper is structured as follows: Section 2 describes the architecture, tasks, and degeneracy metrics. Section 3 presents our empirical findings, and Section 4 discusses their broader implications.
@@ 2.1 Model architecture and training procedure
[…]
@@ 2.2 Task suite for diagnosing solution degeneracy
We use a diverse set of four tasks that elicit different neural dynamics: pattern recognition (N-Bit Flip-Flop), delayed decision-making (Delayed Discrimination), pattern generation (Sine Wave Generation), and evidence accumulation (Path Integration). Below, we briefly describe the task structure and the typical dynamics required to solve each one.
@@
Figure 2: Our task suite spans memory, integration, pattern generation, and decision-making. Each task is designed to place distinct demands on the network’s dynamics. N-Bit Flip-Flop: The network must remember the last nonzero input on each of $N$ independent channels. Delayed Discrimination: The network compares the magnitude of two pulses, separated by a variable delay, and outputs their sign difference. Sine Wave Generation: A static input specifies a target frequency, and the network generates the corresponding sine wave over time. Path Integration: The network integrates velocity inputs to track position in a bounded 2D or 3D arena (schematic shows 2D case).
@@
N-Bit Flip-Flop Task Each RNN receives $N$ independent input channels taking values in $\{-1,0,+1\}$, which switch with probability $p_{\text{switch}}$. The network has $N$ output channels that must retain the most recent nonzero input on their respective channels. The network dynamics form $2^{N}$ fixed points, corresponding to all binary combinations of $\{-1,+1\}^{N}$.
@@
Delayed Discrimination Task The network receives two pulses of amplitudes $f_{1},f_{2}\in[2,10]$, separated by a variable delay $t\in[5,20]$ time steps, and must output $\operatorname{sign}(f_{2}-f_{1})$. In the $N$-channel variant, comparisons are made independently across channels. The network forms task-relevant fixed points to retain the amplitude of $f_{1}$ during the delay period.
@@
Sine Wave Generation The network receives a static input specifying a target frequency $f\in[1,30]$ and must generate the corresponding sine wave $\sin(2\pi ft)$ over time. We define $N_{\text{freq}}$ target frequencies, evenly spaced within the range $[1,30]$, and use them during training. In the $N$-channel variant, each input channel specifies a frequency, and the corresponding output channel generates a sine wave at that frequency. For each frequency, the network dynamics form and traverse a limit cycle that produces the corresponding sine wave.
@@
Path Integration Task Starting from a random position in 2D, the network receives angular direction $\theta$ and speed $v$ at each time step and updates its position estimate. In the 3D variant, the network takes as input azimuth $\theta$, elevation $\phi$, and speed $v$, and outputs updated $(x,y,z)$ position. The network performs path integration by accumulating velocity vectors based on the input directions and speeds. After training, the network forms a Euclidean map of the environment in its internal state space.
@@ 2.3.1 Behavioral degeneracy
We define behavioral degeneracy as the variability in network responses to out-of-distribution (OOD) inputs. We measure OOD performance as the mean squared error of all converged networks that achieved near-asymptotic training loss under a temporal generalization condition. For the Delayed Discrimination task, we doubled the delay period. For all other tasks, we doubled the length of the entire trial to assess generalization under extended temporal contexts. Behavioral degeneracy is defined as standard deviation of the OOD losses: $\sigma_{\mathrm{OOD}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{L}_{\mathrm{OOD}}^{(i)}-\overline{\mathcal{L}}_{\mathrm{OOD}}\right)^{2}}$, where ¯$\overline{\mathcal{L}}_{\mathrm{OOD}}$ is the mean OOD loss.
@@ 2.3.2 Dynamical degeneracy
We use Dynamical Similarity Analysis (DSA) [67] to compare the neural dynamics of task-trained networks through pairwise analyses. DSA compares the topological structure of dynamical systems and has been shown to be more robust to noise and better at identifying behaviorally relevant differences than prior metrics such as Procrustes Analysis and Central Kernel Alignment [35]. For a pair of networks $X$ and $Y$, DSA identifies a linear forward operator for each system—$A_{x}$ and $A_{y}$—which maps neural activity from one time step to the next. These operators are then compared up to a rotation. The DSA distance between two systems is computed by minimizing the Frobenius norm between the operators, up to rotation:
@@
where $O(n)$ is the orthogonal group. We define dynamical degeneracy as the average DSA distance across all network pairs. Additional details are provided in Appendix D. We focus on comparing neural dynamics because RNNs implement computations through time-evolving trajectories rather than static input representations. In addition, we assess representational degeneracy using Singular Vector Canonical Correlation Analysis (SVCCA) [71]. As shown in Appendix E, the four factors that influence dynamical degeneracy do not impose the same constraints on representational degeneracy.
@@ 2.3.3 Weight degeneracy
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the recurrent weight matrices for a pair of RNNs, $\mathcal{P}(n)$ is the set of permutation matrices of size $n\times n$, and $\|\cdot\|_{F}$ denotes the Frobenius norm. See Appendix D.2 for additional details. For comparing networks of different sizes, we normalize the above norm by the number of parameters in the weight matrix.
@@ 3 Results
3.1 Task compleity modulates degeneracy across levels
@@
Figure 3: Higher task complexity reduces dynamical and behavioral degeneracy, but increases weight degeneracy. (A) Two-dimensional MDS embedding of network dynamics shows that independently trained networks converge to more similar trajectories as task complexity increases. (B) Dynamical, (C) weight, and (D) behavioral degeneracy across 50 networks as a function of task complexity. Shaded area indicates ±$\pm 1$ standard error.
At the behavioral level, networks trained on more complex tasks consistently showed lower variability in their responses to OOD test inputs (Figure 3D). This finding suggests that increased task complexity, by reducing dynamical degeneracy, also leads to more consistent—and less degenerate—OOD generalization behavior across networks. Together, the results at the behavioral and dynamical levels support the Contravariance Principle, which posits an inverse relationship between task complexity and the dispersion of network solutions [10].
At the weight level, we found that pairwise distances between converged RNNs’ weight matrices increased consistently with task complexity (Figure 3C). This likely reflects increased dispersion of local minima in weight space for harder tasks. This interpretation is consistent with prior work on mode averaging and loss landscape geometry in feedforward networks. These studies show that harder tasks tend to yield increasingly isolated minima, separated by steeper barriers [33, 25, 54, 22, 1, 70, 55]. A complementary perspective comes from [48], who introduced the intrinsic dimension—the lowest-dimensional weight subspace that still contains a solution—which can serve as a proxy for task complexity. As task complexity increases, the intrinsic dimension of the weight space expands and each solution occupies a thinner slice of a higher-dimensional space, leading to minima that lie further apart. In Section 3.2, we propose an additional mechanism: an interaction between task complexity and the network’s learning regime that further amplifies weight-space degeneracy.

### v3

@@ Abstract
Task-trained recurrent neural networks (RNNs) are widely used in neuroscience and machine learning to model dynamical computations. To gain mechanistic insight into how neural systems solve tasks, prior work often reverse-engineers individual trained networks. However, different RNNs trained on the same task and achieving similar performance can exhibit strikingly different internal solutions, a phenomenon known as solution degeneracy. Here, we develop a unified framework to systematically quantify and control solution degeneracy across three levels: behavior, neural dynamics, and weight space. We apply this framework to 3,400 RNNs trained on four neuroscience-relevant tasks: flip-flop memory, sine wave generation, delayed discrimination, and path integration, while systematically varying task complexity, learning regime, network size, and regularization. We find that higher task complexity and stronger feature learning reduce degeneracy in neural dynamics but increase it in weight space, with mixed effects on behavior. In contrast, larger networks and structural regularization reduce degeneracy at all three levels. These findings empirically validate the Contravariance Principle and provide practical guidance for researchers seeking to tune the variability of RNN solutions, either to uncover shared neural mechanisms or to model the individual variability observed in biological systems. This work provides a principled framework for quantifying and controlling solution degeneracy in task-trained RNNs, offering new tools for building more interpretable and biologically grounded models of neural computation.
@@ 1 Introduction
Recurrent neural networks (RNNs) are widely used in machine learning and computational neuroscience to model dynamical processes. They are typically trained with standard nonconvex optimization methods and have proven useful as surrogate models for generating hypotheses about the neural mechanisms underlying task performance [sussillo2014neural, rajan2016recurrent, barakRecurrentNeuralNetworks2017, mastrogiuseppe2018linking, vyas2020computation, driscoll_flexible_2024]. Traditionally, the study of task-trained RNNs has focused on reverse-engineering a single trained model, implicitly assuming that networks trained on the same task would converge to similar solutions, even when initialized or trained differently. However, recent work has shown that this assumption does not hold universally, and the solution space of task-trained RNNs can be highly degenerate: networks may achieve the same level of training loss, yet differ in out-of-distribution (OOD) behavior, internal representations, neural dynamics, and connectivity [turner2021charting, maheswaranathan2019universality, kurtkaya2025dynamicalphasesshorttermmemory, pagan_individual_2025, clark_symmetries_2025, lappalainen_connectome-constrained_2024, murray2025phasecodesemergerecurrent]. For instance, [maheswaranathan2019universality] found that while trained RNNs may share certain topological features, their representational geometry can vary widely. Similarly, [turner2021charting] showed that task-trained networks can develop qualitatively distinct neural dynamics and OOD generalization behaviors.
These findings raise fundamental questions about the solution space of task-trained RNNs: What factors govern the solution degeneracy across independently trained RNNs? When the solution space of task-trained RNNs is highly degenerate, to what extent can we trust conclusions drawn from a single model instance? While feedforward networks have been extensively studied in terms of how weight initialization and stochastic training (e.g., mini-batch gradients) lead to divergent solutions, RNNs still lack a systematic and unified understanding of the factors that govern solution degeneracy [das2020systematic, turner2024simplicity, martinelli2024expandandclusterparameterrecoveryneural, martinelli2025flatchannelsinfinityneural, fort2019deep, goodfellow2015qualitativelycharacterizingneuralnetwork, li2018visualizinglosslandscapeneural, jastrzębski2018factorsinfluencingminimasgd, chaudhari2017entropysgdbiasinggradientdescent, kornblith2019similarityneuralnetworkrepresentations]. Cao and Yamins [cao2024explanatory] proposed the Contravariance Principle, which posits that as the computational objective (i.e., the task) becomes more complex, the solution space should become less dispersed—since fewer models can simultaneously satisfy the stricter constraints imposed by harder tasks. While this principle is intuitive and compelling, it has thus far remained largely theoretical and has not been directly validated through empirical studies.
@@
In this paper, we introduce a unified framework for quantifying solution degeneracy at three levels: behavior, neural dynamics, and weight space (Figure 1). Leveraging this framework, we isolate four key factors that control solution degeneracy: task complexity, learning regime, network width, and structural regularization. We apply this framework in a large-scale experiment, training 50 independently initialized RNNs on each of four neuroscience-relevant tasks. By systematically varying task complexity, learning regime, network width, and regularization, we map how each factor shapes degeneracy across behavior, dynamics, and weights. We find that as task complexity increases (whether via more input–output channels, higher memory demand, or auxiliary objectives, or as networks undergo stronger feature learning), their neural dynamics become more consistent, while their weight configurations grow more variable. In contrast, increasing network size or imposing structural regularization during training reduces variability at both the dynamics and weight levels. At the behavioral level, each of these factors reliably modulates behavioral degeneracy; however, the relationship between behavioral and dynamical degeneracy is not always consistent.
Table 1 summarizes how task complexity, learning regime, network size, and regularization affect degeneracy across levels. In both machine learning and neuroscience, the desired level of degeneracy may vary depending on the specific research questions being investigated. This framework offers practical guidance for tailoring training to a given goal, whether encouraging consistency across models [kepple2022curriculum], or promoting diversity across learned solutions [liebana_garcia_striatal_2023, fascianelli_neural_2024, PPR:PPR811803].
@@
  • A systematic sweep of four factors: task complexity, feature learning, network size, and regularization, and a summary of their effects across levels (Table 1), with practical guidance for tuning consistency vs. diversity [kepple2022curriculum, liebana_garcia_striatal_2023, fascianelli_neural_2024, PPR:PPR811803].
@@ 2.1 Model architecture and training procedure
We use discrete-time nonlinear vanilla recurrent neural networks (RNNs), defined by the update rule: $\mathbf{h}_{t}=\tanh\left(\mathbf{W}_{h}\mathbf{h}_{t-1}+\mathbf{W}_{x}\mathbf{x}_{t}+\mathbf{b}\right)$ where $\mathbf{h}_{t}\in\mathbb{R}^{n}$ is the hidden state, $\mathbf{x}_{t}\in\mathbb{R}^{m}$ is the input, $\mathbf{W}_{h}\in\mathbb{R}^{n\times n}$ and $\mathbf{W}_{x}\in\mathbb{R}^{n\times m}$ are the recurrent and input weight matrices, and $\mathbf{b}\in\mathbb{R}^{n}$ is a bias vector. A learned linear readout is applied to the hidden state to produce the model’s output at each time step. Networks are trained with Backpropagation Through Time (BPTT) [werbos1990backpropagation], which unrolls the RNN over time to compute gradients at each step. All networks are trained using supervised learning with the Adam optimizer without weight decay. Learning rates are tuned per task (Appendix B). For each task, we train 50 RNNs with 128 hidden units. Weights are initialized from the uniform distribution $\mathcal{U}\left(-1/\sqrt{n},\,1/\sqrt{n}\right)$ and hidden states are initialized to be zeros.
In all experiments, we train networks until them reach a near‑asymptotic, task‑specific mean-squred error (MSE) threshold on the training set (see Appendix B), after which we allow a patience period of 3 epochs and stop training to measure degeneracy. This early‑stopping criterion ensures that networks trained on the same task achieve comparable final losses before any degeneracy analysis.
@@ 2.2 Task suite for diagnosing solution degeneracy
We selected a diverse set of four tasks designed to elicit distinct neural dynamics commonly studied in neuroscience. The N-Bit Flip-Flop task captures pattern recognition and memory retrieval processes, analogous to Hopfield-type attractor networks that store discrete binary patterns and retrieve them from partial cues [hopfield1982neural, jarne2024exploring]. The Delayed Discrimination task models working memory maintenance in classic delayed-response paradigms [funahashi1989mnemonic, goldmanrakic1995cellular]. The Sine Wave Generation task represents pattern generation, analogous to Central Pattern Generators (CPGs) that produce self-sustaining rhythmic outputs underlying motor control [marder2001central], as well as oscillatory activity observed in motor cortex during movement [churchland2012neural]. Finally, the Path Integration task is inspired by hippocampal and entorhinal circuits that build a cognitive map of the environment to track position by integrating self-motion cues [mcnaughton2006path]. These tasks have also been used in prior benchmark suites for neuroscience-relevant RNN training [yang2019task, khona2023winning, maheswaranathan2019universality], underscoring their broad relevance for studying diverse neural computations. Below, we briefly describe the task structure and the typical dynamics required to solve each one.
@@
Figure 2: Our task suite spans memory, integration, pattern generation, and decision-making. Task schematics and representative network trajectories projected onto the top principal components are shown in (A)–(B). The four tasks are: N-Bit Flip-Flop: The network must remember the last nonzero input on each of $N$ independent channels. Delayed Discrimination: The network compares the magnitude of two pulses, separated by a variable delay, and outputs their sign difference. Sine Wave Generation: A static input specifies a target frequency, and the network generates the corresponding sine wave over time. Path Integration: The network integrates velocity inputs to track position in a bounded 2D or 3D arena (schematic shows 2D case).
@@
N-Bit Flip-Flop Task Each RNN receives $N$ independent input channels taking values in $\{-1,0,+1\}$, which switch with probability $p_{\text{switch}}$. The network has $N$ output channels that must retain the most recent nonzero input on their respective channels. The network dynamics form $2^{N}$ fixed points, corresponding to all binary combinations of $\{-1,+1\}^{N}$. The output range of this task is [1,1][-1,1] and we apply an early-stopping training MSE threshold at 0.001.
@@
Delayed Discrimination Task The network receives two pulses of amplitudes $f_{1},f_{2}\in[2,10]$, separated by a variable delay $t\in[5,20]$ time steps, and must output $\operatorname{sign}(f_{2}-f_{1})$. In the $N$-channel variant, comparisons are made independently across channels. The network forms task-relevant fixed points to retain the amplitude of $f_{1}$ during the delay period. The output range of this task is [1,1][-1,1] and we apply an early-stopping training MSE threshold at 0.01.
@@
Sine Wave Generation The network receives a static input specifying a target frequency $f\in[1,30]$ and must generate the corresponding sine wave $\sin(2\pi ft)$ over time. We define $N_{\text{freq}}$ target frequencies, evenly spaced within the range $[1,30]$, and use them during training. In the $N$-channel variant, each input channel specifies a frequency, and the corresponding output channel generates a sine wave at that frequency. For each frequency, the network dynamics form and traverse a limit cycle that produces the corresponding sine wave. The output range of this task is [1,1][-1,1] and we apply an early-stopping training MSE threshold at 0.05.
@@
Path Integration Task Starting from a random position in 2D, the network receives angular direction $\theta$ and speed $v$ at each time step and updates its position estimate. In the 3D variant, the network takes as input azimuth $\theta$, elevation $\phi$, and speed $v$, and outputs updated $(x,y,z)$ position. The network performs path integration by accumulating velocity vectors based on the input directions and speeds. After training, the network forms a map of the environment in its internal state space. The output range of this task is [5,5][-5,5] and we apply an early-stopping training MSE threshold at 0.05.
In our task suite, trained RNNs develop distinct stable dynamical objects: fixed-point (N-Bit Flip Flop, Delayed Discrimination), limit cycle (Sine Wave Generation), and attractor manifold (Path Integration). In Appendix E, we extend our task suite to include a next-step prediction task on the Lorenz 96 chaotic attractors [lorenz1996predictability], where networks exhibit chaotic dynamical regime.
@@ 2.3.1 Behavioral degeneracy
We define a novel metric for behavioral degeneracy as the variability in network responses to out-of-distribution (OOD) inputs. We quantify OOD performance as the mean squared error of all converged networks that achieved near-asymptotic training loss under a temporal generalization condition. For the Delayed Discrimination task, we doubled the delay period. For all other tasks, we doubled the length of the entire trial to assess generalization under extended temporal contexts. Behavioral degeneracy is defined as standard deviation of the OOD losses: $\sigma_{\mathrm{OOD}}=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\mathcal{L}_{\mathrm{OOD}}^{(i)}-\overline{\mathcal{L}}_{\mathrm{OOD}}\right)^{2}}$, where $\overline{\mathcal{L}}_{\mathrm{OOD}}$ is the mean OOD loss. While we focus primarily on the temporal generalization condition for behavioral degeneracy since it directly probes RNNs’ sequence processing capacities and their ability to generalize across extended temporal horizons, the same metric can be readily applied to other OOD conditions, such as input noise or external perturbations. In the rest of the paper, we use the term behavioral degeneracy [temporal generalization] to explicitly indicate the OOD condition being tested.
@@ 2.3.2 Dynamical degeneracy
We use Dynamical Similarity Analysis (DSA) [ostrow_beyond_2023] to compare the neural dynamics of task-trained networks through pairwise analyses. While previous comparison methods mostly focus on geometry of the data [raghu2017svcca, Kriegeskorte2008RSA, williams_generalized_2022, Schrimpf2020BrainScore], RNNs implement computations through time-varying trajectories rather than static representations, and two RNNs exhibiting similar representational geometry can implement distinct dynamical computations, and vise versa. DSA compares the topological structure of the neural dynamics and has been shown to be more robust to noise and better at identifying behaviorally relevant differences than geometry-based comparison method [guilhot2025dynamical]. For a pair of networks $X$ and $Y$, DSA projects their time series of activities to a higher-dimensional space and identifies a linear dynamic operator for each system via next-step prediction. The DSA distance between two systems is then computed by minimizing the Frobenius norm between the operators, up to an orthogonal transformation (rotation and reflection):
@@
where $O(n)$ is the orthogonal group. We define dynamical degeneracy as the average DSA distance across all network pairs. Additional details on the DSA metric are provided in Appendix F. We note that scale of the DSA distance used to quantify dynamical degeneracy can depend on the choice of DSA hyperparameters. To ensure fair comparison across conditions, we keep all DSA hyperparameters fixed for RNNs trained on the same task. To assess if the neural dynamics across different trained networks are statistically different, we also establish a null distribution by comparing neural trajectories sampled from the same underlying network, see Appendix F.3 for details.
We focus on comparing neural dynamics because RNNs implement computations through time-evolving trajectories rather than static input representations. In addition, we assess representational degeneracy using Singular Vector Canonical Correlation Analysis (SVCCA) [raghu2017svcca]. As shown in Appendix G, the four factors that influence dynamical degeneracy do not impose the same constraints on representational degeneracy.
@@ 2.3.3 Weight degeneracy
where $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are the recurrent weight matrices for a pair of RNNs, $\mathcal{P}(n)$ is the set of permutation matrices of size $n\times n$, and $\|\cdot\|_{F}$ denotes the Frobenius distance. See Appendix F.2 for additional details. For comparing $d_{PIF}$ computed on networks of different sizes, we normalize the above norm by the number of parameters in the weight matrix.
@@ 3 Results
3.1 Task complexity modulates degeneracy across levels
@@
Figure 3: Higher task complexity reduces dynamical and behavioral degeneracy, but increases weight degeneracy. (A) Two-dimensional MDS embedding of network dynamics shows that independently trained networks converge to more similar trajectories as task complexity increases. (B) Dynamical, (C) weight, and (D) behavioral degeneracy [temporal generalization] across 50 networks as a function of task complexity. Shaded area indicates $\pm 1$ standard error.
At the behavioral level, networks trained on more complex tasks consistently showed lower variability in their responses to OOD test inputs (Figure 3D) in the temporal generalization condition. This finding suggests that increased task complexity, by reducing dynamical degeneracy, also leads to more consistent and less degenerate behavior on the temporal generalization condition across networks. Together, the results at the behavioral and dynamical levels support the Contravariance Principle, which posits an inverse relationship between task complexity and the dispersion of network solutions [cao2024explanatory].
At the weight level, we found that pairwise distances between converged RNNs’ weight matrices increased consistently with task complexity (Figure 3C). This likely reflects increased dispersion of local minima in weight space for harder tasks. This interpretation is consistent with prior work on mode averaging and loss landscape geometry in feedforward networks, showing that harder tasks tend to yield increasingly isolated minima, separated by steeper barriers [goodfellow2015qualitatively, frankle2020linear, lucas2021monotonic, fort2019largescale, achille2019information, qu2024rebasin, ly2025multifractal]. A complementary perspective comes from [li2018intrinsic] who introduced the intrinsic dimension as the lowest-dimensional weight subspace that still contains a solution, which can serve as a proxy for task complexity. As task complexity increases, the intrinsic dimension of the weight space expands and each solution occupies a thinner slice of a higher-dimensional space, leading to minima that lie further apart. In Section 3.2, we propose an additional mechanism: an interaction between task complexity and the network’s learning regime that further amplifies weight-space degeneracy.